### v2

# Emergent spin order and steady-state superradiance in one-dimensional baths

Abstract
Spontaneous collective decay in incoherently driven atomic ensembles generates macroscopic coherence in the steady state, as exemplified by superradiant lasing in single-mode cavities. Whether spontaneous order persists in multimode reservoirs with competing collective decay channels and light propagation remains an open question. We address this problem by analyzing atoms coupled to one-dimensional electromagnetic baths through two models: a ring cavity with two bright decay channels, and a bidirectional waveguide where, in addition to competition between channels, propagation induces Hamiltonian dipole–dipole interactions. For suitable pumping strengths, both models enter a synchronization window leading to steady-state phase ordering and superradiant emission, scaling as $N^{2}$. The resulting order is not described by a single macroscopic dipole: in the ring cavity spontaneous chirality emerges at the level of individual trajectories, while the waveguide develops a local chirality with different orders dominating opposite ends of the array. The spectral analysis of the emitted light seems to indicate line narrowing with increased system size in the ring cavity. Line narrowing in the waveguide remains inconclusive within accessible numerics. Our results demonstrate how competition and propagation shape emergent order beyond the Dicke limit.
I INTRODUCTION
The spontaneous emergence of order in systems composed of many interacting subsystems is pervasive across biology, chemistry, and the social sciences [27]. In physics, disorder–order transitions have been extensively studied in thermodynamic equilibrium, yet some of the most striking manifestations occur far from equilibrium, where a featureless external energy input can stabilize robust ordered steady states and generate macroscopic coherence. A prime example is the conventional laser [27, 28, 38], where atoms in a “high-Q” single-mode cavity pumped above the lasing threshold, generate coherent light via stimulated emission. The intracavity field both quantifies the degree of organization and provides the feedback that enforces it.
Another example of spontaneous order in the form of macroscopic quantum coherence is that of a superradiant laser, realized in the “bad-cavity” limit where photons leak out faster than the time needed for stimulated emission [42, 41, 73, 36, 50]. Here, the cavity acts as a shared electromagnetic reservoir that mediates all-to-all interactions between the atoms, and endows the system with permutational symmetry if the atoms are placed at the cavity antinodes. As a result, the atoms spontaneously decay via a single bright jump operator [18]. For suitable pumping strengths, the atoms relax to a superradiant state [18], emitting a field whose intensity scales quadratically with atom number and whose linewidth can fall below that of a single-atom, suppressing noise from dephasing or decay outside of the cavity. The repeated action of the single jump operator synchronizes the atomic phases [2, 66, 76], establishing a macroscopic dipole that is resilient against quantum fluctuations. In contrast to a conventional laser, phase information is thus stored primarily in the atomic ensemble rather than in the cavity field.
An open question is whether steady-state spontaneous spin order – together with superradiant light emission and spectral narrowing – can be stabilized away from the previously described “Dicke limit”. If atoms decay into a multimode electromagnetic bath with continuous spectrum, the competition from the resulting collective jump operators – together with light-propagation effects encoded in the Hamiltonian – may preclude the emergence of an ordered phase. Nevertheless, prior work [76, 9] indicates that the Dicke limit is not required to observe order-like signatures and line narrowing. Furthermore, recent work on transient self-organization in extended systems – including superradiant bursts [40, 39, 57, 51, 35, 52], directional photon emission [18, 11, 37], and phase ordering detected via heterodyne measurements [71] – is encouraging. Yet a clear physical picture of how competition and propagation shape the emergent order remains elusive, as does the question of whether linewidths below the natural single-atom value and steady-state superlinear (i.e., superradiant) emission are feasible beyond the Dicke limit.
Here, we address these questions by elucidating the roles of multimode competition and light propagation in the steady-state emission of atoms coupled to one-dimensional (1D) reservoirs. We consider two spin models that progressively incorporate these effects: (i) a ring cavity [34], which supports two bright jump operators, each driving competing orders for the atomic coherence and emitted light, and (ii) a bidirectional waveguide, where in addition to competition between atomic collective modes, light propagation manifests through coherent dipole–dipole interactions mediated by the waveguide. Beyond serving as minimal models to explore competition and propagation, these scenarios are directly relevant to experimental platforms such as cold atoms coupled to microring resonators [75, 60, 74], ring and bow-tie cavities [17, 15, 14] and nanofibers [62, 22, 24, 58].
The spontaneous order arising from the synchronization of atomic phases leads to two very distinct behaviors in ring cavities and waveguides. In a ring cavity, there are two competing macroscopic orders that maximize either left- or right-propagating photon emission. Shot-to-shot a single order dominates, in an example of dynamical spontaneous mirror symmetry breaking. In a waveguide, the order results in a phase-separated state where the right and the left orders coexist, each dominating opposite ends, with a domain wall region in the middle of the array. In both cases, the optimal intensity scales as $N^{2}$, as in a superradiant laser. The thresholds for collective emission are reduced compared to those in a cavity, but remain finite, leaving a sizable synchronization window. We also analyze the spectral properties of the emitted light and find that the linewidth narrows with increasing atom number in the ring cavity, whereas our results on line narrowing in the waveguide are inconclusive for the system sizes explored numerically.
II SPIN MODELS AND METHODS
II.1 Spin models
We consider a system of $N$ two-level atoms, each with a resonance frequency $\omega_{0}$, interacting with a common electromagnetic reservoir at an individual decay rate $\Gamma_{\rm 1D}$. Atoms also undergo parasitic individual decay at rate $\Gamma^{\prime}$ and are incoherently pumped at a rate $w$. After tracing out the electromagnetic field, the atomic dynamics under the Born-Markov approximation are governed by the master equation [25, 20, 5]:
$\dot{\hat{\rho}}=-\frac{{\rm i}}{\hbar}\left[\hat{H},\hat{\rho}\right]+\mathcal{L}_{\text{dis}}[\hat{\rho}]+\mathcal{L}_{\Gamma}[\hat{\rho}]+\mathcal{L}_{w}[\hat{\rho}$], (1)
where (making $\hbar=1$)
$\hat{H}$ $=\sum_{n,m=1}^{N}J_{nm}\hat{\sigma}_{+}^{n}\hat{\sigma}_{-}^{m},$ (2a)
[…]
and the pumping term in the Lindbladian, $\mathcal{L}_{w}[\hat{\rho}$], is obtained from $\mathcal{L}_{\Gamma}[\hat{\rho}$] by swapping $\hat{\sigma}_{+}^{n}\leftrightarrow\hat{\sigma}_{-}^{n}$. Here, $\hat{\sigma}_{+}^{n}\equiv\mathinner{|{e_{n}}\rangle}\mathinner{\langle{g_{n}}|}$ denotes the coherence operator for emitter $n$. The coefficients $\{J_{nm},\Gamma_{nm}\}$ encode the coherent and dissipative interactions, respectively, between atoms $n$ and $m$. These interactions are mediated by the shared environment and are fully determined by the electromagnetic field propagator (i.e., the electromagnetic Green’s function) between pairs of atoms, evaluated at the resonance frequency [5].
Figure 1: Schematics of the spin models for a single-mode cavity, ring cavity and waveguide, and their corresponding spin orders. In all settings, NN two-level atoms with parasitic decay $\Gamma^{\prime}$ and incoherent pumping $w$ are coupled to a shared electromagnetic environment, into which they emit photons at rate $\Gamma_{\text{1D}}$. In a single-mode cavity (a), atoms provide global feedback to one another, leading to a synchronized steady state with a single phase. In a ring cavity (b), atoms provide competing directional feedback, driving the ensemble to synchronize to emit all the light into either the right- or the left-propagating mode. In a bidirectional waveguide (c), two types of feedback are also present. However, feedback is position dependent, leading to a phase-separated steady state.
Collective dissipation, described by $\mathcal{L}_{\text{dis}}[\hat{\rho}$] in the master equation, can be recast in terms of (atomic) collective jump operators by diagonalizing the matrix of coefficients $(\bm{\Gamma})_{nm}\equiv\Gamma_{nm}$ as
[…]
where $\hat{\mathcal{O}}_{\nu}=\sum_{n=1}^{N}b_{\nu}^{n}\hat{\sigma}_{-}^{n}$ are collective jump operators with rates $\Gamma_{\nu}$ (and with spatial profile given by the eigenvectors $\bm{b}_{\nu}$ of $\bm{\Gamma}$). These modes may be superradiant ($\Gamma_{\nu}>\Gamma_{\rm 1D}$) or subradiant ($\Gamma_{\nu}<\Gamma_{\rm 1D}$). Alternatively, $\mathcal{L}_{\text{dis}}[\hat{\rho}$] can be rewritten in terms of directional jump operators [12, 16].
The steady state of the density matrix dynamics described by Eq. (1) is unique. This follows from Evans’ theorem [21, 8], which states that the steady state is unique whenever the dynamical generators of the master equation – the Hamiltonian and all jump operators – span the full operator space. In our case, this condition is met due to the presence of individual parasitic decay and incoherent pumping processes. Furthermore, notice that the choice of incoherent pumping endows Eq. (1) with global $U(1)$ symmetry, that is, invariance under the transformation $\hat{\sigma}_{+}^{n}\rightarrow e^{i\theta}\hat{\sigma}_{+}^{n}$ $\forall$ $n$. As a consequence of these two observations, and in contrast to the case of coherent pumping, the expectation values of all emitter coherences, and hence all phase information, vanish identically in the steady state for any choice of initial state. As we show throughout this work, this information can nonetheless be recovered by examining either single realizations within the ensemble of quantum trajectories or correlation functions of spin operators.
Signatures of order can also be observed at the ensemble-averaged level in the correlation functions of the emitted electromagnetic field. These can be obtained from the atomic evolution via the input-output relations [33, 10, 69], which express the field in terms of atomic operators and the electromagnetic field propagator. The rate of photon emission into the common reservoir, which is proportional to the field intensity, reads
$R=\sum_{\nu}\Gamma_{\nu}\langle\hat{\mathcal{O}}^{\dagger}_{\nu}\hat{\mathcal{O}}_{\nu}\rangle.$ (4)
Examples of order manifestations include transient superradiant bursts when the atoms decay from a fully inverted state [18], or suppressed emission when subradiant states are populated [4]. Collective behavior manifests itself not only in the total emitted intensity, but also in the spectrum of the emitted light, defined for a field operator $\hat{E}^{+}$ as
$S_{\hat{E}}(\omega)\equiv 2\,\text{Re}\!\left[\int_{0}^{\infty}d\tau\,e^{-{\rm i}\omega\tau}\langle\hat{E}^{-}(t_{ss}+\tau)\hat{E}^{+}(t_{ss})\rangle\right],$ (5)
[…]
We now specify $\{J_{nm},\Gamma_{nm}\}$ for three distinct spin models, which physically correspond to atoms interacting via a single-mode cavity, a ring cavity, and a waveguide, as shown in Fig. 1. These setups differ in their boundary conditions for the electromagnetic field, which in turn leads to different interactions between atoms and thus different steady states. We present these models in order of increasing complexity.
Single-mode cavity (single collective jump operator): Here, we set the Hamiltonian to zero and assume permutational symmetry among the spins, leading to
$J_{nm}=0,\hskip 5.69054pt\Gamma_{nm}=\Gamma_{\rm 1D}.$ (6)
This yields a single bright collective jump operator
$\hat{\mathcal{O}}_{C}=\frac{1}{\sqrt{N}}\sum_{n=1}^{N}\hat{\sigma}_{-}^{n},$ (7)
with collective decay rate $N\Gamma_{\rm 1D}$. The positive frequency component of the field operator is directly related to the jump operator and reads [7]
$\hat{E}_{C}^{+}={\rm i}\frac{\Gamma_{\rm 1D}}{2}\sum_{n=1}^{N}\hat{\sigma}_{-}^{n}.$ (8)
This model, depicted in Fig. 1(a), is realized by atoms resonantly coupled to a bad cavity, i.e., a cavity whose photon leakage rate far exceeds any other timescale [7, 41]. This is the conventional model of Dicke superradiance [18] and is, to date, the only one where superradiant lasing has been demonstrated to exist [42, 41, 6, 30].
Ring cavity (two collective jump operators): For this model, we again set the Hamiltonian to zero but now introduce position-dependent dissipative couplings [29, 34]:
$J_{nm}=0,\hskip 5.69054pt\Gamma_{nm}=\Gamma_{\rm 1D}\cos(k|z_{n}-z_{m}|),$ (9)
where $k$ is the wavenumber of the cavity mode and $z_{n}$ are the atom positions. This yields two bright collective jump operators corresponding to left (counterclockwise) and right (clockwise) emission [47]
$\hat{\mathcal{O}}_{L/R}=\frac{1}{\sqrt{N}}\sum_{n=1}^{N}e^{\pm{\rm i}kz_{n}}\hat{\sigma}_{-}^{n},$ (10)
each with a decay rate of $\Gamma_{L/R}=N\Gamma_{\rm 1D}/2$.
This model, depicted in Fig. 1(b), is physically realized by resonantly coupling atoms to a ring cavity and is designed to probe mode competition. Using the input-output relation [10] together with the Green’s function for a ring cavity [29, 34], the left- and right- propagating fields at any position $z$ are
$\hat{E}_{RC,R}^{+}(z)$ $={\rm i}\frac{\Gamma_{\rm 1D}}{4}\sum_{n=1}^{N}e^{{\rm i}k(z-z_{n})}\hat{\sigma}_{-}^{n},$ (11a)
$\hat{E}_{RC,L}^{+}(z)$ $={\rm i}\frac{\Gamma_{\rm 1D}}{4}\sum_{n=1}^{N}e^{{\rm i}k(z_{n}-z)}\hat{\sigma}_{-}^{n},$ (11b)
which are proportional, up to an overall constant factor, to the operators $\hat{\mathcal{O}}_{L/R}$.
Bidirectional waveguide (two jump operators and Hamiltonian evolution): Finally, this model incorporates both Hamiltonian evolution and position-dependent dissipation. The interaction coefficients are given by [5, 47]:
$J_{nm}-{\rm i}\frac{\Gamma_{nm}}{2}=-{\rm i}\frac{\Gamma_{\rm 1D}}{2}e^{{\rm i}k|z_{n}-z_{m}|}.$ (12)
This equation yields dissipative coefficients $\Gamma_{nm}=\Gamma_{\rm 1D}\cos(k|z_{n}-z_{m}|)$, identical to those of the ring cavity [Eq. (9)], but now also introduces a coherent exchange $J_{nm}=(\Gamma_{\rm 1D}/2)\sin(k|z_{n}-z_{m}|)$. Since the dissipative coefficients are the same as in the ring cavity, the collective decay is still described by the left/right modes $\hat{\mathcal{O}}_{L/R}$ [Eq. (10)]. This model, illustrated in Fig. 1(c), is physically realized by atoms coupled to a bidirectional waveguide that supports a guided mode with wavenumber $k$, and serves as a minimal example that incorporates both mode competition and light propagation. By the input-output relation, the right- and left- propagating emitted fields at any position $z$ along the waveguide are
$\hat{E}_{WG,R}^{+}(z)$ $={\rm i}\frac{\Gamma_{\rm 1D}}{2}\sum_{z>z_{n}}e^{{\rm i}k(z-z_{n})}\hat{\sigma}_{-}^{n},$ (13a)
$\hat{E}_{WG,L}^{+}(z)$ $={\rm i}\frac{\Gamma_{\rm 1D}}{2}\sum_{z<z_{n}}e^{{\rm i}k(z_{n}-z)}\hat{\sigma}_{-}^{n}.$ (13b)
When the observation point $z$ lies far to the right (left) of all the atoms, the expression for $\hat{E}_{WG,R}^{+}(z)$ ($\hat{E}_{WG,L}^{+}(z)$) reduce, up to a global phase, to the right (or left) jump operator of Eq. (10).
For both the ring cavity and the waveguide, we assume that atoms form ordered arrays with spacing $d$, i.e., $z_{n}=nd$, although as we will detail later, the results are expected to extend to disordered arrays. In the “mirror” configuration ($kd=n\pi$ with $n\in\mathbb{Z}$), the coherent part $J_{nm}$ vanishes, and the system supports a single bright mode, reducing the dynamics to that of a single-mode cavity [13].
II.2 Analytical and computational methods
Ordered phases of matter are emergent phenomena that become more robust and well-defined as $N$ grows, which makes the computation of their physical properties challenging away from thermal equilibrium. Equation (1) describes an open quantum system with an exponentially large Hilbert space, rendering exact naive simulations intractable at large $N$. When full permutation symmetry is present (as in a single-mode cavity) the dynamics can be computed exactly with computational cost $\sim N^{3}$ [67]. This can be extended to ring cavities by exploiting partial permutation symmetry in commensurate “superspin” arrays [34]. In particular, for spacings $d=m\lambda_{0}/(2p)$ with $m,p\in\mathbb{Z}$, the ensemble decomposes into $p$ symmetric groups $\mathcal{S}_{\alpha}$ ($\alpha=1,\dots,p$), each represented by a collective spin
$\hat{J}^{\alpha}_{-}=$ $\frac{1}{N_{\alpha}}\sum_{n\in\mathcal{S}_{\alpha}}c_{n,\alpha}\hat{\sigma}_{-}^{n},$ (14a)
$\hat{J}^{\alpha}_{z}=$ $\frac{1}{2N_{\alpha}}\sum_{n\in\mathcal{S}_{\alpha}}\hat{\sigma}_{z}^{n},$ (14b)
where $N_{\alpha}$ is the number of atoms in $\mathcal{S}_{\alpha}$ and the coefficients $\{c_{n,\alpha}\}$ ensure that coherences of atoms within the same symmetric group add in phase (see Appendix A). This reduces the computational complexity to $\sim N^{3p}$ and enables a compact description fully in terms of the collective operators in Eqs. (14). The case $p=1$ reduces to the single macroscopic dipole describing a superradiant laser [67]. Advancing beyond the single-mode analysis of Ref. [41], we employ this formalism in Sec. IV to characterize the steady-state emission of the ring cavity.
In the waveguide, where both full and partial permutation symmetry are absent, we resort to approximate methods. The simplest is mean field, which neglects correlations by factorizing the density matrix and treating each emitter as a classical dipole. To include leading quantum corrections, we employ the Truncated Wigner Approximation (TWA) for dissipative spin systems, which yields $2N$ stochastic differential equations, where correlated white noise terms model quantum fluctuations [49, 55, 44, 43, 26]. The reliability of this approach is supported by benchmarks against exact small-$N$ simulations, together with recent evidence that few-body observables in superradiant decay with all-to-all coupling are well captured by mean-field trajectories under heterodyne unraveling [72]. We use this method throughout the manuscript to assess the impact of quantum fluctuations.
II.3 Mean field and classical synchronization
Employing a mean-field approximation provides analytical insight into the emergent spin order by interpreting it as phase synchronization. Here, we follow Ref. [76] and present mean-field equations that we will later apply to our particular spin models. In the mean-field approach, quantum correlations are neglected by assuming a factorized density matrix, $\rho(t)\approx\bigotimes_{n=1}^{N}\rho_{n}(t)$. Each atom is then described by an inversion, coherence, and a phase, $\{s_{n}^{z},s_{n}^{\perp},\phi_{n}\}$, defined as
$2s^{z}_{n}$ $\equiv\langle\hat{\sigma}_{z}^{n}\rangle,$ (15a)
$s^{\perp}_{n}e^{i\phi_{n}}$ $\equiv\langle\hat{\sigma}_{+}^{n}\rangle,$ (15b)
which reduce Eq. (1) to the mean-field equations [76]
$\dot{s}^{z}_{\ell}$ $=-s^{\perp}_{\ell}\sum_{n\neq\ell}s^{\perp}_{n}\left(2J_{n\ell}\sin(\phi_{n\ell})+\Gamma_{n\ell}\cos(\phi_{n\ell})\right)$
$-(\Gamma^{\prime}+\Gamma_{1D}+w)s^{z}_{\ell}+\frac{w-\Gamma^{\prime}-\Gamma_{1D}}{2},$ (16a)
$\dot{s^{\perp}_{\ell}}$ $=s^{z}_{\ell}\sum_{n\neq\ell}s^{\perp}_{n}\left(2J_{\ell n}\sin(\phi_{n\ell})+\Gamma_{\ell n}\cos(\phi_{n\ell})\right)$
$-\frac{\Gamma^{\prime}+w+\Gamma_{1D}}{2}s^{\perp}_{\ell},$ (16b)
$\dot{\phi}_{\ell}$ $=\frac{s^{z}_{\ell}}{s^{\perp}_{\ell}}\sum_{n\neq\ell}s^{\perp}_{n}\left(-2J_{\ell n}\cos(\phi_{n\ell})+\Gamma_{\ell n}\sin(\phi_{n\ell})\right),$ (16c)
where $\phi_{n\ell}\equiv\phi_{n}-\phi_{\ell}$. Relative phases determine the coherent and collective dissipative evolution, as seen from the terms proportional to $J_{nm}$ and $\Gamma_{nm}$. In all mean-field simulations, we take the atoms to be fully excited at initial time and incorporate fluctuations by sampling initial conditions from the distribution $p(s^{z},s^{\perp},\phi)\propto e^{-2{s^{\perp}}^{2}}\delta\left(\sqrt{{\left(s^{\perp}\right)}^{2}+{\left(s^{z}\right)}^{2}}-\tfrac{1}{2}\right)$ [55].
Within mean field, phase dynamics map to a Sakaguchi–Kuramoto model [76], a standard description of synchronization and pattern formation in networks of classical limit cycles [2]. The basic Kuramoto model for phases $\{\phi_{\ell}\}$ assumes permutational symmetric coupling and reads [32, 59, 2]
$\dot{\phi}_{\ell}=\delta_{\ell}+K\sum_{\ell}^{N}\sin(\phi_{n\ell}),$ (17)
where $K>0$ is a coupling constant and $\delta_{\ell}$ the natural frequency of oscillator $\ell$. Synchronization is captured by the order parameter
$Z=\sum_{\ell}^{N}e^{{\rm i}\phi_{\ell}}\equiv re^{i\psi},$ (18)
whose magnitude $r$ quantifies phase alignment. In terms of $r$ and $\psi$, the phase dynamics becomes
$\dot{\phi}_{\ell}=\delta_{\ell}+Kr\sin(\psi-\phi_{\ell}),$ (19)
showing that phases are attracted toward the order parameter phase $\psi$, which causes the growth of $r$, increasing attraction and establishing a positive feedback loop that stabilizes at nonzero $r$. Cosine-type phase interaction (like the ones describing the coherent interactions in the phase mean-field equation) give instead an integrable model [63]: its evolution retains memory of the initial conditions, and disordered configurations do not spontaneously synchronize. When sine- and cosine-type all-to-all couplings compete, the balance between the two dictates whether the system synchronizes or disorder persists [46]. Extensions beyond permutation-symmetric coupling can yield reduced synchronization or chimera states, where synchronized and unsynchronized oscillators coexist [53, 31, 1, 46].
Even without coherent interactions, Eq. (16c) yields richer behavior than the standard Kuramoto model [76]. Effective couplings are inhomogeneous and time dependent through $s_{\ell}^{z}$ and $s_{\ell}^{\perp}$. Positive steady-state couplings require $s_{\ell}^{z}>0$ and $s^{\perp}_{\ell}\neq 0$, which define a finite synchronization window for $w$. When $w\in[w^{(l)},w^{(u)}]$, the couplings are positive at long times and phase synchronization is possible. The boundaries of the synchronization window are set by $\{J_{nm},\Gamma_{nm}\}$.
The concept of using a macroscopic order parameter, as is done in the Kuramoto model [Eq. (19)], extends naturally to our spin models. In what follows, we focus on the emergence of order within the synchronization window, at the mean-field level and including low-order fluctuations via TWA. A quantitative analysis of the synchronization thresholds and properties of the emitted field is deferred to Sec. IV.
III EMERGENT SPIN ORDER
We begin by reviewing a single-mode bad cavity, the simplest setting for atomic synchronization [42, 41, 66, 76]. We then add layers of complexity via one-dimensional reservoirs: a ring cavity with two competing bright jump operators, and a bidirectional waveguide where competition coexists with propagation encoded by coherent interactions.
III.1 Single-mode cavity: superradiant lasing
Figure 2: Order in the single-mode cavity. Wigner function for the electromagnetic field emitted by $N=1024$ atoms coupled to a single-mode cavity, for $\Gamma^{\prime}=20\Gamma_{\rm 1D}$ and $w=N\Gamma_{\rm 1D}/2$. White squares indicate examples of mean-field results.
For this spin model – defined by the couplings in Eq. (6) – the expectation value of the field is the order parameter. The (unnormalized) mean-field approximation for the cavity field in Eq. (8) reads
$E=\sum_{n=1}^{N}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv re^{-{\rm i}\psi},$ (20)
where $r$ and $\psi$ denote the field amplitude and phase. According to Eq. (16c), phases have an all-to-all homogeneous coupling [Fig. 1(a)] as in the Kuramoto model. In terms of $r$ and $\psi$, Eq. (16c) becomes
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{s^{\perp}_{\ell}}r\sin(\psi-\phi_{\ell}).$ (21)
For $w$ within the synchronization window this realizes a Kuramoto model with positive coupling, leading to phase synchronization [76]. In the steady state, atoms emit a field with fixed amplitude $r$ but an arbitrary global phase $\psi$, reflecting the global $U(1)$ symmetry of the master equation. As in conventional single-mode lasers, quantum fluctuations induce phase diffusion of the mean-field solution [38, 61]. In the complex $E$-plane, the Wigner function forms a ring with radius set by $r$, indicating fixed intensity but undefined phase, as illustrated in Fig. 2. Shot-to-shot the phase is picked by spontaneous symmetry breaking.
III.2 Ring cavity: winner takes all
Figure 3: Order in the ring cavity. Mean-field trajectories and Wigner function in (a) the space of right- and left-field amplitudes, and (b) in the complex plane of the right field. Diamonds indicate the final configurations of the mean-field dynamics. Time is encoded in the line transparency (fainter for earlier times), and each trajectory is colored according to which field dominates in the steady state. In (b), the Wigner function in regions far from the center have been multiplied by a factor of 10 for visualization purposes. Panel (c) shows the magnetization $\mathcal{M}$ for a single realization of the phase dynamics, where the right field prevails. Parameters: $N=1024$, $\Gamma^{\prime}=0$, $w=N\Gamma_{\rm 1D}/4$.
In the ring cavity, the competition between the two modes leads to dynamical spontaneous symmetry breaking of both $U(1)$ (global phase) and mirror (left vs right order) symmetries at the level of individual realizations, as we demonstrate below. In analogy with the single-mode cavity case, we define classical order parameters for left and right emission with amplitudes $r_{L/R}$ and phases $\psi_{L/R}$ as
$E_{L/R}=\sum_{n=1}^{N}e^{\pm ikz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{L/R}e^{-{\rm i}\psi_{L/R}}.$ (22)
These fields allow us to write the phase dynamics of Eq. (16c) compactly as
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{2s^{\perp}_{\ell}}\sum_{\nu=\text{L,R}}r_{\nu}\sin(\psi_{\nu}-\theta_{\nu,\ell}-\phi_{\ell}),$ (23)
where we have defined the position-dependent phases $\theta_{L/R,n}=\mp kz_{n}$. When pumping produces positive inversion and nonzero coherences, the dynamics involve two competing feedback loops that pull $\phi_{\ell}$ toward either $\phi_{\ell}^{R}=\psi_{R}-kz_{n}$ or $\phi_{\ell}^{L}=\psi_{L}+kz_{n}$. Each of these attractors maximizes either $r_{R}$ or $r_{L}$, and the strength of attraction grows with the corresponding field amplitude. Thus, the strongest field at a given time attracts phases toward its fixed point, further amplifying itself and suppressing the rival mode. As captured by Eq. (22) and sketched in Fig. 1(b), the light emitted by each atom into the right/left mode serves as feedback to synchronize the rest of the atoms to the right/left order.
Within the synchronization window, as shown in Fig. 3(a) and (b), mean-field trajectories relax to one of two steady states: either the left field acquires a finite amplitude while the right field is suppressed, or vice versa, reminiscent of a conventional ring-laser behavior [38]. To visualize the microscopic phase order, we define the “magnetization” for atom $n$ at time $t$ as
$\mathcal{M}(n,t)=\frac{|\Delta\phi_{n}(t)-kd|-|\Delta\phi_{n}(t)+kd|}{2kd},$ (24)
where $\Delta\phi_{n}\equiv\phi_{n}-\phi_{n-1}$. The magnetization quantifies whether neighbors are closer to maximizing right ($\mathcal{M}(n,t)\rightarrow 1$) or left ($\mathcal{M}(n,t)\rightarrow-1$) emission. Figure 3(c) shows that, after an initial transient, all the phases stabilize to either the right or the left order.
Low-order quantum fluctuations give rise to a bimodal Wigner distribution concentrated near the two mean-field solutions. In the complex plane of either field, the distribution features both a peak at the origin and an outer ring. These correspond, respectively, to realizations in which the right-field mode is suppressed and the left is stabilized at a finite amplitude with an undetermined phase (on average), or vice versa.
As discussed in Sec. II, the steady state of Eq. (1) is unique. The two solutions therefore do not correspond to distinct steady states of the master equation. Individual decay and incoherent pumping reduce the mirror-inversion symmetry of Eq. (1) for the ring-cavity model to a weak symmetry [8, 3] (i.e., a symmetry preserved at the level of the master equation but not at the level of single trajectories). Consistent with this, the two peaks in the Wigner distribution result from stochastic switching between the two dynamical outcomes in measurement-conditioned trajectories [64], in contrast to dissipative freezing that occurs in the presence of a strong symmetry [54]. As the atom number increases, the Wigner distribution becomes increasingly localized around a single mean-field solution. Consequently, any individual trajectory or experimental realization will remain localized near one of the two mean-field solutions, effectively breaking the symmetry.
The locking to either a global left or right phase order motivates an ansatz for the steady state as an incoherent superposition of spin-coherent states with phases matching the two possible orders [72, 71]. We harness the form of the ansatz recently proposed in Ref. [71] for transient superradiant dynamics in a waveguide (but here employed to describe the steady state in a ring cavity), which reads
$\hat{\rho}_{\text{ans,RC}}(t_{ss})=\frac{1}{4\pi}\int_{0}^{2\pi}d\psi\left[\hat{\rho}_{R}(\psi)+\hat{\rho}_{L}(\psi)\right],$ (25)
where
$\begin{aligned} \hat{\rho}_{L/R}(\psi)&=\bigotimes_{n=1}^{N}\mathinner{|{\theta_{n},\pm kz_{n}+\psi}\rangle}\mathinner{\langle{\theta_{n},\pm kz_{n}+\psi}|},\end{aligned}$ (26)
and spin coherent states are defined as $\mathinner{|{\theta,\phi}\rangle}=\cos\frac{\theta}{2}\mathinner{|{e}\rangle}+e^{{\rm i}\phi}\sin\frac{\theta}{2}\mathinner{|{g}\rangle}$. The free angle $\psi$ in Eq. (25) reflects the global $U(1)$ symmetry. Heterodyne unraveling [64] is particularly well-suited to observe the ordering encoded in Eq. (25): although the steady state of Eq. (1) has vanishing atomic coherences due to the $U(1)$ symmetry, measurement backaction generates coherence at the level of individual trajectories. As a result, each realization develops nonzero coherences and, after sufficiently long evolution, tends to synchronize to either the left or the right order.
Observables sensitive to the synchronization direction average to zero in the full density matrix, since roughly half of the realizations lock to each side. Nonetheless, signatures of the order are observable at the level of correlation functions [76]
$C_{nm}=\langle\hat{\sigma}_{+}^{n}\hat{\sigma}_{-}^{m}\rangle(t=t_{ss}).$ (27)
Evaluating the above equation using $\hat{\rho}_{\text{ans,RC}}(t_{ss})$, we find that interference between equal weights of left- and right-ordered realizations produces a cosine spatial pattern
$C_{nm}$ $\propto\frac{1}{2}\left(e^{-{\rm i}k(z_{n}-z_{m})}+e^{{\rm i}k(z_{n}-z_{m})}\right)$ (28)
$=\cos(k(z_{n}-z_{m})),$
which in agreement with the behavior observed in the TWA simulations (not shown).
III.3 Waveguide: coexistence of left and right order
Figure 4: Order in a bidirectional waveguide. (a) Mean-field magnetization $\mathcal{M}$ for a single realization of the phase dynamics with parameters $(N,\Gamma^{\prime},w,kd)=(1024,0,N\Gamma_{\rm 1D}/8,2\pi/3)$. (b) Schematics of the correlation phase matrix $\arg[C_{nm}]$ together with representative plots. Labels (I)–(V) denote regions with distinct behavior, as discussed in the main text. The phase correlations $\arg[C_{nm}]$ along the horizontal cuts shaded in black in the schematics, $\arg[C_{1m}]\simeq kd(1-m)$ and $\arg[C_{Nm}]\simeq-kd(N-m)$, are shown in the upper and lower panels on the right. These correlation phases align with the values predicted by the left and right ordered states as $m$ moves deeper into the corresponding domain. Correlation phases along the diagonals $(n,n+d)$ are plotted as a function of the rescaled and shifted coordinate $n^{\prime}$ defined in the main text (middle right panel). Plots in (b) are done with parameters $(N,\Gamma^{\prime},w,kd)=(200,2\Gamma_{\rm 1D},N\Gamma_{\rm 1D}/4,2\pi/3)$. TWA simulation results are shown in colors and ansatz predictions in black.
In addition to the competition between jump operators, the waveguide has a non-zero Hamiltonian that encodes the propagation of light [23], dramatically altering the steady state. Heuristically, the distinction from a ring cavity lies in boundary conditions. In a ring, light emitted by any atom into either mode loops around to provide global feedback to the entire ensemble [see the definition of the order parameters in Eq. (22)]. In a waveguide the feedback is directional, as light emitted by atom $\ell$ to the right mode only serves as feedback for atoms located to its right, as sketched in Fig. 1(c). In the mean-field treatment, this asymmetry suggests the definition of right and left position-dependent order parameters
$E_{R}(\ell)$ $=\sum_{z_{n}<z_{\ell}}e^{-{\rm i}kz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{R}(\ell)e^{-{\rm i}\psi_{R}(\ell)},$ (29a)
$E_{L}(\ell)$ $=\sum_{z_{n}>z_{\ell}}e^{{\rm i}kz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{L}(\ell)e^{-{\rm i}\psi_{L}(\ell)},$ (29b)
a construction with parallels to earlier approaches that introduced position-dependent order parameters to describe synchronization in systems with inhomogeneous couplings and frustration [31, 1]. In terms of these, the equations for atoms coupled to a waveguide mirror those of the ring cavity, except that the order parameters are now position-dependent. For example, the equation that describes the phase evolution reads
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{s^{\perp}_{\ell}}\sum_{\nu=\text{L,R}}r_{\nu}(\ell)\sin(\psi_{\nu}(\ell)-\theta_{\nu,\ell}-\phi_{\ell}).$ (30)
In the mean-field treatment, each phase thus evolves under position–dependent feedback. Emission into the right mode accumulates along the array, peaking at the right edge, where atoms are coherent and lock their phases to radiate to the right. Symmetrically, atoms on the left lock to the left order. The impact of coherent interactions is made evident by the magnetization $\mathcal{M}(n,t)$, shown in Fig. 4(a) for a trajectory of the classical dynamics. Unlike the ring cavity [Fig. 3(c)], where a single macroscopic order dominates, atoms in a bidirectional waveguide undergo phase separation between the two orders. Most realizations of the mean field dynamics thus reach a steady state where the right and left field intensity are approximately identical.
This phase–separated order can be probed beyond mean field by analyzing the steady-state coherence matrix $C_{nm}$ (defined in Eq. (27) and computed via TWA), and particularly its phase, $\text{arg}[C_{nm}]$. Unlike in the ring cavity, atoms at the edges consistently lock to either the left or the right mode, making the mean-field phase pattern directly visible in the steady state whenever both atoms belong to the same domain [see Fig. 4(b)]. In contrast, correlations between atoms in different domains gradually deviate from this ordered phase structure.
To understand the physical origin of these correlations, we introduce a phenomenological ansatz that captures their essential features. As in the ring cavity, the emergent order motivates an ansatz as an incoherent superposition of spin-coherent components. Here, however, each realization of the steady state displays coexisting left and right orderings. Therefore, the ansatz for the ring cavity in Eq. (25) is no longer applicable. Instead, we propose
$\hat{\rho}_{\text{ans,WG}}(t_{ss})=\int_{0}^{2\pi}\frac{d\psi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\psi_{L}}{2\pi}\sum_{\textbf{A}\in\{R,L\}^{N}}P(\textbf{A})$ (31)
$\times\bigotimes_{n=1}^{N}\mathinner{|{\theta_{n},\phi(A_{n},n,\psi_{A_{n}})}\rangle}\mathinner{\langle{\theta_{n},\phi(A_{n},n,\psi_{A_{n}})}|},$
with
$\phi(A_{n},n,\psi_{A_{n}})=\begin{cases}kz_{n}+\psi_{L}&\text{ if }A_{n}=L\\ -kz_{n}+\psi_{R}&\text{ if }A_{n}=R.\end{cases}$ (32)
Here, $P(\mathbf{A})$ denotes the probability of a given configuration of local orderings. For example, $\mathbf{A}=[L,L,R,L\ldots]$ means that the first two atoms align with the left order, the third aligns with the right order, and so on. The independent integration over $\psi_{R}$ and $\psi_{L}$ reflects the assumption that the relative phase between the left and right ordered domains is completely undetermined in the steady state (at least, on average). It also respects the global $U(1)$ symmetry.
To fully specify the ansatz, one must provide the configuration probabilities $P(\mathbf{A})$. However, to compute the correlation matrix elements $C_{nm}$, only the marginals $P(A_{n},A_{m})$ are required. We propose a specific form for $P(A_{n},A_{m})$ that incorporates the fact that atoms on the left (right) side of the array preferentially synchronize to the left (right) order, while central atoms choose either with comparable probability. As shown in Fig. 4(b), this model successfully reproduces the characteristic structure of the correlation matrix $C_{nm}$. Details on the specific probability model, fitting parameters and the correlation calculation are given in Appendix B.
Equipped with the ansatz in Eq. (31), we are now in a position to provide physical intuition for the behavior of the correlations $C_{nm}$ in the distinct regions indicated in Fig. 4(b):
  • “Phase-locked” regions [(I) and (III)]: These correspond to correlations between pairs of atoms that both belong to either the left or the right domain. In these regions, the phase of the correlation follows $\arg[C_{nm}]\simeq k(z_{n}-z_{m})$ for the left domain and $\arg[C_{nm}]\simeq-k(z_{n}-z_{m})$ for the right domain.
  • “Domain wall” region [(II)]: This region involves correlations between atoms in the central portion of the array. Atoms near the middle synchronize with the left or right order with roughly equal probability. Consequently, as in the ring cavity [Eq. (28)], the average correlations here follow $C_{nm}\simeq\cos(k(z_{n}-z_{m}))$. Additionally, right- and left- propagating fields in this section of the array are weaker than in the borders, so local feedback is reduced. As a result, these atoms are more susceptible to fluctuations and desynchronize more easily.
  • Interdomain correlations [(IV) and (V)]: These correspond to atom pairs located in opposite domains. For a fixed relative phase between orders, each pair exhibits a well-defined phase relation, $\arg[C_{nm}]\simeq\pm[k(z_{n}+z_{m})+\psi_{L}-\psi_{R}]$. However, averaging over $\psi_{R}$ and $\psi_{L}$ washes out this phase relation, restoring a cosine-like pattern.
To visualize these features, we plot the phase of the correlations along the horizontal and diagonal cuts indicated in Fig. 4(b). Each diagonal is labeled by an integer $d\in[-(N-1),N-1]$, with $|d|=1$ corresponding to nearest-neighbor correlations $C_{(n+1)n}$, $|d|=2$ to next-nearest neighbors $C_{(n+2)n}$, and so on. All cuts are compiled in Fig. 4(b), where we collapse the correlations of all diagonals and show $\arg[C_{(n+d)n}]$ as a function of the shifted and rescaled coordinate $n^{\prime}=\frac{2(n-1)+d}{2(N-1)}\in[0,1]$. Under this mapping, correlations between atoms in the left domain (region (I)) are sent to $n^{\prime}\sim 0$, those in the right domain (region (III)) to $n^{\prime}\sim 1$, while correlations involving atoms across domains or near the center of the array (regions (II), (IV) and (V)) accumulate around $n^{\prime}\sim 0.5$. Correlations near $n^{\prime}=0$ and $1$ follow the expected left- and right-ordered phases, whereas a smooth crossover appears around $n^{\prime}=0.5$, where the correlation phase clusters around 0 or $\pi$, consistent with the cosine-like behavior discussed above.
We note that the ansatz in Eq. (31) is a phenomenological model intended to reproduce the behavior of correlations after averaging over trajectories. A more detailed analysis of individual trajectories may provide further insight on how quantum fluctuations influence the mean-field order. For example, although the relative phase between the left and right order parameters is assumed to be arbitrary, more general models could be developed by introducing a probability distribution of the relative phase, $P(\psi_{R}-\psi_{L})$. The present ansatz reproduces the correct average correlations provided that the probability distribution of this relative phase is an even function. A possible refinement of the model would be to allow for atoms that are not synchronized with either order parameter. Nonetheless, the minimal model in Eq. (31) already demonstrates that the mean-field intuition is consistent with the TWA numerical results.
Finally, we remark that for both the ring cavity and the waveguide, the mean-field analysis does not require the atoms to form an ordered array. We therefore expect similar phase-ordered states to emerge even in disordered ensembles.
IV PROPERTIES OF THE EMITTED LIGHT
Here we determine the pumping thresholds for synchronization and superradiant emission, and show that the emitted intensity scales as $N^{2}$ for all three spin models. Using an extension of the TWA to compute multi-time correlators [26] together with an $SU(4)$ representation of the Liouvillian for the single-mode cavity [65], we then analyze the degree of first-order coherence of the field by studying the spectral linewidth.
IV.1 Thresholds and optimal intensity
Figure 5: Thresholds for superradiant light emission and scaling of maximum emission rate. (a) Emission rate and (b) excited-state population in the steady state as a function of pumping strength. Insets: zoom on the first threshold. (c) Maximum intensity vs atom number. In all plots, solid lines show the analytical expressions [from Ref. [41] for the single-mode cavity and from Eqs. (33) for the ring cavity], points show the numerical results with TWA, and the dashed line in (c) indicates the estimate of $R_{\text{max}}$ for the waveguide described in the main text. Parameters: $(N,\Gamma^{\prime},kd)=(1024,2\Gamma_{\rm 1D},2\pi/3)$.
In a cavity, the lower and upper thresholds for superradiant lasing are known to be $w^{(l)}=\Gamma^{\prime}$ and $w^{(u)}=N\Gamma_{\rm 1D}$, respectively, with a maximum emission rate scaling as $R_{\text{max}}=N^{2}\Gamma_{\rm 1D}/8$ [42, 41]. The thresholds are intuitively clear since incoherent pumping must exceed the independent decay rate to generate positive inversion and enable synchronization, while excessively strong pumping destroys correlations seeded by collective decay. Since $N\Gamma_{\rm 1D}$ sets the maximum collective decay timescale, for $w>w^{(u)}$ correlations are washed out faster than they can form.
We find analytical expressions for the ring-cavity thresholds for superradiant emission by exploiting the superspin method introduced in Ref. [34] and described in Sec. II. Analyzing the Heisenberg evolution of the collective operators in Eq. (14) and performing a second-order cumulant expansion analogous to the single-mode case (a summary of this calculation is given in Appendix A), we find that the ring-cavity steady state also exhibits threshold behavior. In particular, the steady-state total intensity and excited–state population $P_{e}\equiv\sum_{n}\langle\hat{\sigma}_{ee}^{n}\rangle$ read
$R(t=t_{ss})$ $=\begin{cases}0\text{ if }w\notin[\Gamma^{\prime},N\Gamma_{\rm 1D}/2]\\ \begin{aligned} \frac{wp}{2N_{s}\Gamma_{\rm 1D}}\left(1-\frac{\Gamma^{\prime}}{w}-\frac{2}{w}\frac{(w+\Gamma^{\prime})^{2}}{N\Gamma_{\rm 1D}}\right)\\[6.99997pt] \text{ if }w\in[\Gamma^{\prime},N\Gamma_{\rm 1D}/2],\end{aligned}\end{cases}$ (33a)
$P_{e}(t=t_{ss})$ $=\begin{cases}\frac{(w-\Gamma^{\prime})p}{2(w+\Gamma^{\prime})}+\frac{1}{2}\text{ if }w\notin[\Gamma^{\prime},N\Gamma_{\rm 1D}/2]\\ \frac{(w+\Gamma^{\prime})p}{N\Gamma_{\rm 1D}}+\frac{1}{2}\text{ if }w\in[\Gamma^{\prime},N\Gamma_{\rm 1D}/2].\end{cases}$ (33b)
Figures 5(a,b) show good agreement between these expressions and the values obtained with TWA. Numerical results for non-commensurate configurations with $d\neq m\lambda_{0}/2p$ follow the same behavior, indicating that the steady-state values of these observables are largely insensitive to the spatial arrangement of the atoms. Equations (33) predict sharp changes in population and the onset and quenching of steady-state collective emission at a lower threshold $w^{(\ell)}=\Gamma^{\prime}$ and an upper threshold $w^{(u)}=N\Gamma_{\text{1D}}/2$. The insets of Fig. 5(a,b) zoom into the behavior around the first threshold. Below $w^{(\ell)}$, TWA yields a small unphysical negative decay rate, consistent with the fact that the steady state in this regime is known to be dark and highly entangled for a single-mode cavity [56], a scenario expected to carry over to 1D reservoirs. Despite this limitation, TWA correctly captures the location of $w^{(\ell)}$. Moreover, Eq. (33a) predicts a maximum emission rate at $w=N\Gamma_{\text{1D}}/4$, with $R_{\text{max}}=N^{2}\Gamma_{\text{1D}}/16$, in agreement with numerical results as shown in Fig. 5(c). As for a single-mode cavity, the upper threshold reflects the requirement that the pump strength remain below the maximum dissipation rate of the system [41, 45], which also sets the fastest synchronization timescale in Eq. (23).
For the waveguide, we obtain the thresholds numerically. These results are in good agreement with a simple back-of-the-envelope analytical calculation, described below. We find $w^{(\ell)}\sim\Gamma^{\prime}$ for all array spacings, but the exact position of the upper threshold depends on lattice constant. Away from the “mirror” configuration, we estimate the upper threshold from Eqs. (29). Because the feedback is position-dependent, each atom has a different maximum pumping at which it can remain synchronized. For instance, atom 1 experiences feedback from (N1)(N-1) atoms driving it toward the left order, so the maximum pumping strength for synchronization is $w^{(u)}(n=1)\sim(N/2-1)\Gamma_{\rm 1D}$. More generally, an atom in the first half of the array will synchronize only if $w<w^{(u)}(n)\sim\left(N/2-n\right)\Gamma_{\rm 1D}$. Averaging over the array then yields $\bar{w}^{(u)}\sim N\Gamma_{\rm 1D}/4$, and by similar reasoning, an optimal pumping $\bar{w}^{(opt)}\sim N\Gamma_{\rm 1D}/8$. Counting atoms with $w^{(u)}(n)<\bar{w}^{(opt)}$ that synchronize to either order at the optimal pump (denoted $N_{\text{sync}}$) and using the known single-mode cavity scaling, we estimate the waveguide intensity at optimal pumping to be $R_{\text{max}}\sim 2\left(N_{\text{sync}}^{2}\Gamma_{\rm 1D}/8\right)=9N^{2}\Gamma_{\rm 1D}/256$. These estimates qualitatively agree with the numerical results in Fig. 5. In contrast to transient superradiance, where coherent interactions can be often neglected when analyzing the superradiant burst [39, 11], the coherent interactions intrinsic to the waveguide strongly modifies the properties of the radiation emitted in the steady state.
IV.2 Spectral properties
Figure 6: Linewidth of the emitted field. (a) Linewidth vs. pumping strength for $N=50$ atoms for the single-mode cavity, ring cavity and waveguide models. (b) Scaling of the linewidth vs NN. In both panels, points are obtained using the TWA, dashed lines show exact results based on the $SU(4)$ formalism, and solid lines serve as guides to the eye. For the ring cavity and waveguide, $(kd,\Gamma^{\prime})=(2\pi/3,2\Gamma_{\rm 1D})$.
In a single-mode cavity, the emitted light is coherent and there is a superradiant “lasing” regime if the pumping is between thresholds [42]. This means that the order emerging from synchronization in a single-mode cavity is robust against fluctuations, allowing steady-state correlations to persist for long times [42]. The two-time correlation of any pair of atoms decays exponentially with the time delay $\tau$, i.e.,
$\langle\hat{\sigma}_{-}^{n}(t_{ss}+\tau)\hat{\sigma}_{+}^{m}(t_{ss})\rangle\sim e^{-\tau/t_{\text{coh}}},$ (34)
with a coherence time $t_{\text{coh}}$ that increases with $N$ [41]. The enhanced lifetime is directly reflected in the spectrum of the light emitted by the cavity, given for large atom number by
$S_{\hat{E}^{+}_{C}}(\omega)\simeq N(N-1)\,\text{Re}\!\left[\int_{0}^{\infty}d\tau\,e^{-{\rm i}\omega\tau}\langle\hat{\sigma}_{-}^{1}(t_{ss}+\tau)\hat{\sigma}_{+}^{2}(t_{ss})\rangle\right],$
which exhibits a linewidth much narrower than that of independent atoms, $(\Delta\nu)_{\text{indep}}=w+\Gamma^{\prime}+\Gamma_{\rm 1D}$. In the large-$N$ limit, the linewidth approaches $(\Delta\nu)_{\text{min}}\sim\Gamma_{\rm 1D}$, so correlations can persist for a maximum time $t_{\text{coh}}\sim 1/\Gamma_{\rm 1D}$ even when parasitic decay dominates over cavity coupling, $\Gamma^{\prime}\gg\Gamma_{\rm 1D}$ [42]. The residual broadening is primarily due to phase fluctuations of the emitted field [68]: as shown in Fig. 2, the amplitude is fixed while the phase remains undetermined. Reduced phase noise enables such light to be used to lock a laser, or as an active frequency reference.
As discussed in Sec. III, synchronization beyond the single-cavity similarly generates phase correlations across the atomic array, but it remains unclear whether these emerging orders are also capable to radiate light with phase robustness that is strengthened as $N$ grows. To address this question, we analyze steady-state correlations of the right- and left-emitted fields. Figure 6 summarizes the linewidth behavior as a function of pump strength and atom number for the three spin models. For moderate $N$, the linewidth exhibits a non-monotonic dependence on the pump rate within the synchronization window [Fig. 6(a)], revealing an optimal pump rate that minimizes the linewidth. Linewidths are computed using TWA [26] and validated against exact $SU(4)$ calculations for the cavity, which reproduce the same qualitative trends.
Figure 6(b) shows the minimum linewidth as a function of atom number. Both the spectral linewidth of $\hat{E}^{+}_{C}$ in the single-mode cavity and that of $\hat{E}^{+}_{RC,R/L}$ in the ring cavity exhibit a seemingly downward trend with increasing $N$. In the waveguide, we calculate the linewidth of the right (left) field at an observation point to the right (left) of all atoms. After accounting for the estimation uncertainty (see Appendix C for details), our numerical results do not conclusively demonstrate linewidth narrowing. This inconclusiveness may arise from numerical limitations. For instance, comparison with exact linewidths in the cavity case, obtained using an $SU(4)$ representation of the dynamics [67], shows that the TWA captures the true linewidth narrowing only qualitatively. In addition, narrowing may only emerge at atom numbers beyond our current simulation capabilities. Alternatively, the absence of clear narrowing may be a genuine physical effect, indicating that the spin order in the waveguide is not robust enough to eliminate noise from the parasitic decay.
The seemingly downward trend of the linewidth for the ring cavity is consistent with the picture developed in Sec. III, where the atoms synchronize to the order set by either $\hat{\mathcal{O}}_{R}$ or $\hat{\mathcal{O}}_{L}$. Although these operators no longer correspond to a single macroscopic dipole, they still represent a global phase order that becomes increasingly robust as more atoms participate, since collective feedback suppresses phase fluctuations of individual atoms. The resulting long-lived phase memory is thus directly visible in the far-field spectra. Nonetheless, we remark that true line narrowing (i.e., $\Delta\nu<\Gamma^{\prime}+\Gamma_{\rm 1D}$) is not achieved within the system sizes accessible to our numerics, although both the observed trends and the nature of the synchronized order strongly suggest that such narrowing should occur in the large-$N$ limit. This indicates that superradiant lasing should be observable in a ring cavity, despite mode competition.
V SUMMARY AND OUTLOOK
We have studied steady-state superradiance of incoherently pumped atoms coupled to 1D reservoirs, focusing on two setups: a ring cavity, where atoms collectively decay into left- and right-propagating modes, and a bidirectional waveguide, where coherent exchange processes accompany collective decay. In the ring cavity, individual realizations lock to one of two macroscopic orders that maximize emission into a single direction, whereas in the waveguide coherent interactions enforce coexistence of left and right order, with edge atoms synchronizing to opposite modes. This self-organization shapes the emitted light. For the ring cavity we derived analytical thresholds and optimal intensities, and observed a downward trend in the linewidth of the emitted light. For the waveguide we obtained numerical estimates of thresholds and intensities, but no clear line narrowing was observable within the system sizes available to our numerical methods.
Our results suggest several directions for future work. On the experimental side, they motivate studies of steady-state superradiance in ring resonators, bow-tie cavities, and nanofibers. On the theory side, a more exhaustive analysis of the linewidth in the waveguide setup is needed, for instance through cumulant expansions. Beyond this, it would be valuable to characterize shot-to-shot fluctuations in the ordering through semiclassical simulations of individual trajectories [70, 72] to clarify the stability of the phase-ordered state and the origin, or possible absence, of linewidth narrowing. It would be interesting to determine whether any correlator becomes increasingly robust to fluctuations as $N$ grows.
Finally, a natural extension of this work is to explore the prospects for superradiant lasing in free space, as originally proposed by Dicke [19]. In particular, it would be interesting to see whether collective decay in the absence of coherent interactions generically leads to basins of attraction associated with the brightest collective modes, as in the ring cavity, and whether the coherent interactions generate crossovers between competing directional orders, as observed in the waveguide. Recent work [9] has shown that a few atoms in a 1D array can emit light with a linewidth that converges to the natural single-atom value by limiting pump-induced broadening and driving only a subset of atoms. In practice, however, 1D atomic arrays with ultranarrow transitions may not be suitable to generate light for atomic clock interrogation, since their maximum collective decay rate scales only linearly with $N$ [45], preventing the realization of superradiant lasers (as the emission is too weak even for large systems). Nevertheless, it would be worth investigating whether selectively pumping only part of a higher-dimensional array could overcome this limitation and enable superradiant lasing in free-space geometries.
Acknowledgments – We are grateful to Joseph T. Lee, Ricardo Gutierrez-Jauregui, Xueyue Zhang and Tingyao Wang for discussions. We acknowledge support by the National Science Foundation through the CAREER Award (No. 2047380), the Air Force Office of Scientific Research through their Young Investigator Program (grant No. 21RT0751), the David and Lucile Packard Foundation, and the Chu Family Foundation.
Appendix A LARGE SPIN LIMIT AND THRESHOLDS
Here we provide details of the analytic calculation of steady-state properties and thresholds for the ring cavity, Eqs. (33).
The evolution of any spin operator $\hat{\mathcal{O}}$ follows the Heisenberg equation
$\frac{d\hat{\mathcal{O}}}{dt}=\frac{{\rm i}}{\hbar}\left[\hat{H},\hat{\mathcal{O}}\right]+\mathcal{L}^{\dagger}_{\text{dis}}[\hat{\mathcal{O}}]+\mathcal{L}^{\dagger}_{\Gamma}[\hat{\mathcal{O}}]+\mathcal{L}^{\dagger}_{w}[\hat{\mathcal{O}}]+\mathcal{F}\left[\hat{\mathcal{O}}\right$], (36)
where the adjoints of the dissipators are
[…]
Here $\hat{H}$ is the Hamiltonian defined in Eq. (2a), and $\mathcal{F}\left[\hat{\mathcal{O}}\right$] is quantum Langevin noise, which averages to zero and can be neglected when computing expectation values of $\hat{\mathcal{O}}$ [52].
As mentioned in the main text, for atoms coupled to a ring cavity, spacings satisfying $d=m\lambda_{0}/2p$, with $m,p\in\mathbb{Z}$ exhibit partial permutational symmetry. We therefore represent each permutationally symmetric group $\mathcal{S}_{\alpha}$ ($\alpha=1,\dots,p$) by macroscopic angular-momentum–like operators [34]
$\hat{J}^{\alpha}_{-}=$ $\frac{1}{N_{\alpha}}\sum_{n\in\mathcal{S}_{\alpha}}(-1)^{m(c_{n,\alpha}+1)}\hat{\sigma}_{-}^{n},$ (38a)
$\hat{J}^{\alpha}_{z}=$ $\frac{1}{2N_{\alpha}}\sum_{n\in\mathcal{S}_{\alpha}}\hat{\sigma}_{z}^{n},$ (38b)
where $c_{n,\alpha}=\lfloor(n-1)/\alpha\rfloor+1$ counts how many wavelengths atom $n$ is away from the first atom in its symmetry group, so that the factor $(-1)^{m(c_{n,\alpha}+1)}$ adds contributions within superspin $\mathcal{S}_{\alpha}$ in phase. Notice that because of the $1/N_{\alpha}$ normalization these operators do not form a $SU(2)$ algebra.
Evaluating Eq. (36) for these operators and applying a second-order cumulant expansion [48, 52], the dynamics of $p$ incoherently driven superspins reads
$\frac{d}{dt}\langle\hat{J}_{z}^{\alpha}\rangle$ $=-\sum_{\mu\neq\alpha}\frac{\Gamma_{\alpha\mu}N_{\mu}}{2}\left(\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\mu}\rangle+\langle\hat{J}_{+}^{\mu}\hat{J}_{-}^{\alpha}\rangle\right)$
$-\Gamma_{\rm 1D}N_{\alpha}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\alpha}\rangle-(\Gamma^{\prime}+w)\langle\hat{J}_{z}^{\alpha}\rangle+\frac{w-\Gamma^{\prime}}{2},$ (39a)
$\frac{d}{dt}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\alpha}\rangle$ $=\langle\hat{J}_{z}^{\alpha}\rangle\bigg[2\Gamma_{\rm 1D}N_{\alpha}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\alpha}\rangle+\sum_{\mu\neq\alpha}\Gamma_{\alpha\mu}N_{\mu}\big(\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\mu}\rangle$
$+\langle\hat{J}_{+}^{\mu}\hat{J}_{-}^{\alpha}\rangle\big)\bigg]-(w+\Gamma^{\prime})\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\alpha}\rangle+\frac{w}{N_{\alpha}},$ (39b)
$\frac{d}{dt}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\xi}\rangle$ $=\Gamma_{\rm 1D}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\xi}\rangle\left(N_{\alpha}\langle\hat{J}_{z}^{\alpha}\rangle+N_{\xi}\langle\hat{J}_{z}^{\xi}\rangle\right)$
$+\Gamma_{\xi\alpha}\left(N_{\xi}\langle\hat{J}_{+}^{\xi}\hat{J}_{-}^{\xi}\rangle\langle\hat{J}_{z}^{\alpha}\rangle+N_{\alpha}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\alpha}\rangle\langle\hat{J}_{z}^{\xi}\rangle\right)$
$+\sum_{\mu\neq\alpha,\xi}N_{\mu}\left(\Gamma_{\mu\alpha}\langle\hat{J}_{+}^{\mu}\hat{J}_{-}^{\xi}\rangle\langle\hat{J}_{z}^{\alpha}\rangle+\Gamma_{\xi\mu}\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\mu}\rangle\langle\hat{J}_{z}^{\xi}\rangle\right)$
$-(\Gamma^{\prime}+w)\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\xi}\rangle.$ (39c)
For $p=1$ and large $N$, Eqs. (39) reduce to the superradiant laser equations of Ref. [41], i.e.,
$\frac{d}{dt}\langle\hat{J}_{z}\rangle$ $=-\Gamma_{\rm 1D}N\langle\hat{J}_{+}\hat{J}_{-}\rangle-(\Gamma^{\prime}+w)\langle\hat{J}_{z}\rangle+\frac{w-\Gamma^{\prime}}{2},$ (40a)
$\frac{d}{dt}\langle\hat{J}_{+}\hat{J}_{-}\rangle$ $=-(\Gamma^{\prime}+w)\langle\hat{J}_{+}\hat{J}_{-}\rangle+2N\Gamma_{\rm 1D}\langle\hat{J}_{z}\rangle\langle\hat{J}_{+}\hat{J}_{-}\rangle$. (40b)
The steady-state population and emission rate are in that case
$R(t=t_{ss})$ $=\begin{cases}0\text{ if }w\notin[\Gamma^{\prime},N\Gamma_{\rm 1D}]\\ \begin{aligned} \frac{w}{2N\Gamma_{\rm 1D}}\left(1-\frac{\Gamma^{\prime}}{w}-\frac{1}{w}\frac{(w+\Gamma^{\prime})^{2}}{N\Gamma_{\rm 1D}}\right)\\[6.99997pt] \text{ if }w\in[\Gamma^{\prime},N\Gamma_{\rm 1D}].\end{aligned}\end{cases}$ (41a)
$P_{e}(t=t_{ss})$ $=\begin{cases}\frac{(w-\Gamma^{\prime})}{2(w+\Gamma^{\prime})}+\frac{1}{2}\text{ if }w\notin[\Gamma^{\prime},N\Gamma_{\rm 1D}/2]\\ \frac{(w+\Gamma^{\prime})}{2N\Gamma_{\rm 1D}}+\frac{1}{2}\text{ if }w\in[\Gamma^{\prime},N\Gamma_{\rm 1D}/2].\end{cases}$ (41b)
[…]
  1. 1.
    Since the steady state is unique, single-superspin variables are identical across $\alpha$ in the long time limit regardless of the initial state, i.e. $\langle\hat{J}_{z}^{\alpha}\rangle\equiv\langle\hat{J}_{z}^{1}\rangle$ and $\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\alpha}\rangle\equiv\langle\hat{J}_{+}^{1}\hat{J}_{-}^{1}\rangle$ $\forall$ $\alpha$.
  2. 2.
    Two-superspin correlations depend only on separation. In fact, for all $t$ and $\alpha\neq\beta$, $\langle\hat{J}_{+}^{\alpha}\hat{J}_{-}^{\beta}\rangle(t)=\Gamma_{\alpha\beta}C(t)$, i.e., correlations between different superspins are described by a global function weighted by the coupling.
Harnessing these observations and the identity
$\sum_{\mu=1}^{p}\Gamma_{\alpha\mu}\Gamma_{\mu\xi}=\frac{p}{2}\Gamma_{\rm 1D}\Gamma_{\alpha\xi},$ (42)
the system (39) reduces from $p+p^{2}$ equations to two by defining $R_{\alpha}=\sum_{\mu}\frac{\Gamma_{\mu\alpha}}{\Gamma_{\rm 1D}}\langle\hat{J}_{+}^{\mu}\hat{J}_{-}^{\alpha}\rangle$:
$\frac{d}{dt}\langle\hat{J}_{z}^{\alpha}\rangle$ $=-\Gamma_{\rm 1D}N_{s}R_{\alpha}-(\Gamma^{\prime}+w)\langle\hat{J}_{z}^{\alpha}\rangle+\frac{w-\Gamma^{\prime}}{2},$ (43a)
$\frac{d}{dt}R_{\alpha}$ $=-(\Gamma^{\prime}+w)R_{\alpha}+N\Gamma_{\rm 1D}\langle\hat{J}_{z}^{\alpha}\rangle R_{\alpha}.$ (43b)
Equations (43) are thus the ring-cavity analogue of (40), differing only by the missing factor of two in the nonlinear term for $R_{\alpha}$ and by the replacement $N\to N_{s}$ in the inversion equation. The resulting steady state corresponds to Eqs. (33) in the main text.
Appendix B STEADY-STATE ANSATZ FOR THE WAVEGUIDE
Figure 7: Modeling the emergent phase patterns in the waveguide. (a) Probability distribution $P_{n}(L)$ in Eq. (44) for the $n$-th atom locking to the left, shown for different values of $\beta$ (setting $\alpha=5$) and $\alpha$ (setting $\beta=1$). (b) Correlation phases $\arg[C_{(n+d)n}]=\arg[\mathinner{\langle{\hat{\sigma}_{+}^{n+d}\hat{\sigma}_{-}^{n}}\rangle}]$ plotted against the rescaled variable $n^{\prime}$ described in the main text. We compare numerical TWA results (colors) to the correlations predicted by the ansatz (black) using the fitted parameters $(\alpha,\beta)=(2.492,0.722)$. The example shown corresponds to an array of $N=300$ atoms with lattice constant $kd=\pi/3$, independent decay $\Gamma^{\prime}=2\Gamma_{\rm 1D}$, and pumping $w=50\Gamma_{\rm 1D}$. (c) Fitted values of the ansatz parameters $\alpha$ and $\beta$ for several system sizes and pumping rates, with fixed lattice constant $kd=2\pi/3$ and independent decay $\Gamma^{\prime}=2\Gamma_{\rm 1D}$. For every NN, each point corresponds to a different pumping value $w$, and its opacity is proportional to the decay rate of that configuration relative to the maximal decay rate, as described in Sec. IV.1.
Here, we provide additional details on the steady-state ansatz in Eq. (31) of the main text, denoted as $\hat{\rho}_{\text{ans,WG}}(t_{ss})$. In particular, we describe the model used for the marginals $P(A_{n},A_{m})$ that enter the evaluation of the correlations $C_{nm}$. Guided by mean-field observations that atoms on the right (left) side of the array tend to synchronize with the right (left) propagating mode, we model the probabilities that atom $n$ and $m$ synchronize to the same order as
$P_{nm}(L,L)=\gamma P_{n}(L)P_{m}(L),$ (44a)
$P_{nm}(R,R)=\gamma[1-P_{n}(L)][1-P_{m}(L)],$ (44b)
with
$P_{n}(L)=\frac{1}{2}+\beta\frac{\tanh\left(\alpha\left(\frac{1}{2}-\frac{n-1}{N-1}\right)\right)}{2\tanh(\alpha/2)},\\$ (45)
where $\beta\in[0,1]$ and $\alpha>0$ are parameters that control, respectively, how strongly atoms near the edges are pinned to their corresponding orders and how sharp the crossover is between the two domains near the array center. For values $\beta\approx 0$, the joint probability of two atoms is independent of their relative position, and the phase pattern of $\text{arg}[C_{nm}]$ obtained from $\hat{\rho}_{\text{ans,WG}}(t_{ss})$ matches the behavior characteristic of a ring cavity. In the opposite limit $\beta\approx 1$, the atoms at the edges are fully locked to the corresponding order. Sample plots of this model for different parameters $(\alpha,\beta)$ are shown in Fig. 7(a).
A direct calculation shows that the ansatz leads to a correlation function of the form
$C_{nm}\propto P_{nm}(R,R)e^{{\rm i}k(z_{n}-z_{m})}+P_{nm}(L,L)e^{-{\rm i}k(z_{n}-z_{m})}$ (46)
$+P_{nm}(R,L)e^{-{\rm i}k(z_{n}+z_{m})}\int_{0}^{2\pi}\frac{d\psi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\psi_{L}}{2\pi}e^{{\rm i}(\psi_{R}-\psi_{L})}$
$+P_{nm}(L,R)e^{{\rm i}k(z_{n}+z_{m})}\int_{0}^{2\pi}\frac{d\psi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\psi_{L}}{2\pi}e^{{\rm i}(\psi_{L}-\psi_{R})}.$
The last two terms vanish due to the arbitrary relative phase between the two orders. Therefore, specifying the probabilities $P_{nm}(L,R)$ and $P_{nm}(R,L)$ is not required to determine the correlations. Their contribution is included indirectly through the parameter $\gamma\in[0,1]$, which reduces the probability that atoms $n$ and $m$ lock simultaneously to the same order. This suppresses $P_{nm}(L,L)$ and $P_{nm}(R,R)$ relative to the disordered configurations $A_{n}=L,A_{m}=R]$ and $A_{n}=R,A_{m}=L]$.
The specific parameters of the model $(\alpha,\beta)$ are obtained by fitting the ansatz to the correlation phases extracted from the TWA numerical simulations. As shown in Fig. 4(b) of the main text, and further illustrated for a different parameter set in Fig. 7(b), the steady-state correlations are well reproduced by the probability model in Eq. (44) [in the plot shown in the main text the fitted parameters are $(\alpha,\beta)=(2.162,0.651)$]. The fitted parameters for multiple system configurations are summarized in Fig. 7(b). They do not exhibit a strong dependence on the pumping strength. As $w$ approaches the optimal pumping rate $R_{max}$, the main effect is a reduction in the noise of the correlation phases $\arg[C_{(n+d)n}]$. Notably, the fitted values of $\beta$ tend to saturate as $N$ increases, approaching a value $\beta(N,\bar{w}^{(opt)})\rightarrow\beta^{*}<1$ [Fig. 7(c)]. This indicates that atoms near the edges do not become completely locked to their preferred order, and that some residual frustration persists even in very large arrays, with $N\sim 1000$.
More sophisticated models for the joint probability distribution may be formulated to more accurately reproduce some of the features revealed by mean-field and TWA approaches, such as a site-dependent parameter $\gamma_{nm}$ or alternative functional models for the individual distributions $P_{n}(L)$. Nevertheless, the minimal model presented here already captures the relevant physical behavior, including other configurations of the atomic positions, as shown in Fig  7(c).
Appendix C LINEWIDTH ANALYSIS
Here, we provide further details on the calculation of the linewidth, the scaling of the minimum linewidth with the atom number $N$, and the estimation of error bars. We also discuss the main limitations of the numerical techniques employed.
In the absence of symmetries, the numerical complexity for computing the linewidth of any field using the TWA scales as $\mathcal{O}(MN^{3})$ [26], where $M$ is the number of trajectories required for convergence. To estimate the linewidths shown in Fig. 6(a) of the main text, we compute the two-time correlator of either the cavity field or the right-propagating field and fit it to an exponential decay of the form
$\langle\hat{E}^{-}(t_{ss}+\tau)\hat{E}^{+}(t_{ss})\rangle\sim e^{-\tau\Delta\nu/2}$ (47)
Although fluctuations inherent to the TWA induce a noise floor at late times, preventing the correlator from vanishing completely as $\tau\rightarrow\infty$, we find that the fitting procedure reliably captures the initial exponential decay of the correlator. The error bars in Fig. 6(a) arise from the stochastic nature of the TWA simulations and correspond to three standard deviations estimated from the fitting procedure.
To determine the minimum linewidth, these calculations are repeated for several values of the incoherent pumping rate $w$. To reduce computational cost and mitigate the impact of noisy or outlier realizations, we do not perform a dense scan in $w$. Instead, we evaluate the linewidth at 10 to 20 selected pumping strengths, which are not necessarily uniformly spaced and may not include the value of $w$ that yields the absolute minimum linewidth. To estimate the minimum from these discrete samples, we fit the data to a LogSumExp function, $\frac{1}{p_{0}}\log\left(\exp{(p_{0}(p_{1}w+p_{2}))}+\exp{(p_{0}(p_{3}w+p_{4}))}\right)$, from which we obtain the minimum linewidth. Alternative fitting models, including polynomials and other non-linear functions, can be used, resulting in similar estimates for the minima. The error bars in Fig. 6(b) likewise represent three standard deviations.
The application of these techniques to our data suggests that the approach is generally valid; however, the results reveal inherent noise that hinders robust conclusions. Because of the various fitting procedures and underlying assumptions involved, the method could be refined at different levels of the calculations, making further numerical studies necessary to assess the validity of these methods and to extend our results to larger system sizes.